2- A welding operation will take place on carbon steel. The desired welding speed is around 0.8 in./s. If an arc welding power supply is used with a voltage of 12 V, what current is needed if the weld width is to be 0.2 in? Assume that the weld cross-section is triangular. (20 marks) 3- For question 2, what would be the temperature increase of the entire structure due to the heat input from welding only? Hint: H()DVCA(20
Answer #1

Solution :-

Power of Weld = 12 * current   Watts

Power required to weld = rho * C * ( Tmelt - Tambient) * A * weld speed

                                   = 7860 * 460 * ( 1723 - 298 ) * 1.29 * 10^-5 * 0.02032      (0.8 in /s = 0.02032 m/s )

                                   = 1350.5437   watts

current = 1350.5437 / 12 = 112.545 Amperes